### v1

# Globally Rigid Convex Braced Polygons

Abstract
Determining the global rigidity of a bar framework is known to be a difficult problem. Here we propose a class of frameworks in the plane, braced polygons, that may be globally rigid and are analogous to convex polytopes in $3$-space that are rigid by Cauchy’s Theorem in 1813. We show that all the strictly convex configurations of a given graph, corresponding to a braced polygon (see Section 2), are infinitesimally rigid if and only if they are all globally rigid in the plane in the class of convex braced polygons, Theorem 2.1. Furthermore, we provide techniques that, for many graphs, prove that all strictly convex configurations of the graph of a braced polygon are globally rigid in the plane, and other techniques to find instances when there is a strictly convex braced polygon that is not globally rigid in the plane. As a test case, we use those techniques to determine when a braced polygon is globally rigid or not, when there are $7$ vertices, with $12$ edges in total, and we know that the underlying graph is globally rigid at a generic configuration. Figures 12,14,15. When there are more than $7$ vertices, it looks like global rigidity starts to become difficult again in many cases, but our techniques still work in many cases.
1 Introduction
1.1 History
One of the most fundamental breakthroughs in understanding the rigidity of structures is Cauchy’s 1813 result that a convex triangulated surface is “rigid”. (See [Gruber and Wills(1993)], vol. A 223–271.) This is despite a mistake in his proof of what is now called his “arm lemma”, and this is despite a lack of interest in the subject of rigidity for many years. Nevertheless, the ideas in his proof are correct, and with the passage of time, interest in the subject of rigidity has increased. (See “Rigidity through a projective lens” in [Nixon et al.(2021)Nixon, Schulze, and Whiteley].) The statement of Cauchy’s theorem is clear. There is only one way to assemble a strictly convex triangulated polyhedron, up to congruent copies, that is convex and where vertices are connected with the given correspondence. As a consequence, there is no continuous motion of the polyhedron, where each (triangular) face remains congruent during the motion. The polyhedron is (locally) rigid.
In 1916 Max Dehn [Dehn(1916)] proved that the same strictly triangulated convex polyhedron is infinitesimally rigid in $\mathbb{R}^{3}$. This means that, when there are vectors ${\bf p}^{\prime}=(p^{\prime}_{1},\dots,p^{\prime}_{n})$ attached to the corresponding vertices of the configuration $(G,{\bf p})$ that act as the first derivative of a finite motion, preserving edge (bar) lengths, then they form the derivative of a congruence. This is equivalent to saying that the rank of a matrix, called the rigidity matrix $R({\bf p})$, is maximal, namely $63n-$ 6, where $n$ is the number of vertices of the polyhedron. This also implies that there is no continuous motion, a flex, of the framework that preserves the length of each edge. In the case of a triangulated sphere, each (triangular) face remains congruent during the motion. A priori, Dehn’s theorem looks to be somewhat weaker than Cauchy’s theorem, and, indeed, in 1975 Herman Gluck [Gluck(1975)] showed that there is a relatively easy way to prove Dehn’s infinitesimal theorem using the combinatorial idea of Cauchy’s theorem. It turns out, anyway, that it is easy to show that Cauchy’s theorem directly implies Dehn’s theorem using Proposition 2.2, which is related to some averaging techniques going back to [Connelly and Whiteley(2010)], [Saliola and Whiteley(2004)], and feels like something Minkowski would have done.
Meanwhile, the question of determining the “rigidity” of a bar framework, even in the plane, was considered as a computational question. In this non-necessarily convex setting instead, it was assumed that the coordinates of the configuration ${\bf p}$ are generic, that is, they do not satisfy any non-zero polynomial equation with integer coordinates. So then the (local) rigidity of the framework $(G,{\bf p})$ only depends on the graph $G$ and not on the particular configuration in $\mathbb{R}^{d}$. It was then shown (Geiringer, Laman)[Pollaczek-Geiringer(1926), Laman(1970)] that there is an algorithm to determine generic rigidity for the graph $G$ in $\mathbb{R}^{2}$, and Lovász and Yemini [Lovász and Yemini(1982)], using matriod methods showed that there is a polynomial-time algorithm to compute the generic rigidity of a graph purely combinatorially. Indeed, Bruce Hendrickson and Jacobs [Jacobs and Hendrickson(1997)] have an elementary algorithm, the pebble game, that is easily implemented to compute the generic rigidity of a graph $G$ in the plane.
In $\mathbb{R}^{3}$ and higher, nobody has found a polynomial-time deterministic algorithm to determine generic rigidity of the graph $G$. On the other hand, one can simply compute the rank of the rigidity matrix $R({\bf p})$ for $(G,{\bf p})$ for a randomly chosen configuration ${\bf p}$ in $\mathbb{R}^{d}$, and with high probability that will determine the generic rigidity of the graph $G$.
With Cauchy’s result about strict convexity, there is no need to assume that the configuration is generic. The convexity assumption is enough. On the other hand, Gluck noticed that since the space of strictly convex realizations is an open set, it automatically implies that the graph of a triangulated sphere is generically rigid in $\mathbb{R}^{3}$. (See also the result of Fogelsanger [Fogelsanger(1988)] about the generic rigidity of other triangulated manifolds.)
Another, closely related notion of local rigidity, and infinitesimal rigidity, is the global rigidity of a bar framework $(G,{\bf p})$ in Euclidean space $\mathbb{R}^{d}$. This means that $(G,{\bf p})$ is the only configuration in $\mathbb{R}^{d}$ with the corresponding bars the same length, up to congruent copies. Global rigidity can be much harder to calculate, even when the configuration appears to be generic. But if we are only interested in determining the global rigidity of a framework $(G,{\bf p})$ at generic configurations, things become easier.
Still there is another problem, even if the configuration ${\bf p}$ is known exactly. A test for generic global rigidity (GGR) is that there is an equilibrium stress $\omega$ for a given framework $(G,{\bf p})$ with a corresponding stress matrix $\Omega$ that has maximal rank. The top two frameworks of Figure 1 both have a one-dimensional stress with a stress matrix $\Omega$ that has maximal rank. In this case the rank is $n-d-1=8-2-1=5$, where $n=8$ is the number of vertices, and $d=2$ is the dimension of the ambient space. The frameworks are both infinitesimally rigid in the plane, and the vertices form strictly convex polygons. The two configurations have corresponding edge lengths the same and are not congruent, even though they testify to the generic rigidity of the corresponding graph. When one knows that the graph $G$ is generically globally rigid, in practice, it can be extremely tedious, and at least difficult, to determine the global rigidity of a given framework $(G,{\bf p})$, even in the plane. (In [Hendrickson(1992)] it is shown that this is an NP-complete problem to determine.) It is possible, with a given graph $G$, to find frameworks that are globally rigid in all higher dimensions, starting in a given lower dimension, universal rigidity, as in [Oba and Tanigawa(2023)], but this may not deal with one’s favorite configuration. One could say that when we say that a graph $G$ is generically globally rigid in a given dimension, we do not know what we are talking about.
1.2 Our results
If one is given a particular tensegrity, or a class of tensegrities, as described below, how can one determine global rigidity? We propose, here, some tools to do that.
In Section 2, we define the class of strictly “convex braced polygons”, which are an open set of configurations which correspond to a particular class of graphs. This makes it easier to test if a particular configuration in that class has a strictly convex configuration that is not globally rigid, given that one can know when it is infinitesimally rigidity. For general graphs, it might be more helpful to use something like the “pure condition” in [White and Whiteley(1983)] to find the critical configurations, where the rigidity matrix $R({\bf p})$ drops rank, for instance. See also [Nixon et al.(2021)Nixon, Schulze, and Whiteley] for a history of some of the classical methods for computing the (infinitesimal) rigidity of bar frameworks.
In Section 3, we show that if a 3-connected braced polygon circuit (in the sense matroid theory) is minimal with respect to the braces, then it always has a strictly convex configuration, where the stress is positve on the outer polygon and negative on all the braces, Theorem 3.6. This result is equivalent to an earlier result of [Geleji and Jordán(2013)] using what they call the “unique interval property” (Definition A.1). Our proof here makes no mention of that property. The result implies global rigidity by a theorem of Connelly[Connelly and Guest(2022), Theorem 5.14.4].
In Section 4, we give some examples that cannot be resolved by the methods in the previous sections, as well as the techniques we can use to determine their global rigidity as well as an averaging technique to determine, in many cases, their non-global rigidity. One very useful method is to use superposition of configurations that can work very efficiently with both the stress matrix $\Omega$ and the rigidity matrix $R({\bf p})$. These methods can solve all cases of 7, or fewer, vertices. See Table 14. We also provide a table of the $7$ vertex strictly convex braced polygons that are unique in the class of strictly convex braced polygons, Figure 15, as mentioned above, in the spirit of Cauchy’s Theorem. In some cases, one can do calculations simply by seeing how configurations cross thresholds without doing computer calculations. See Figure 21 for an example, and Remark 4.1 for an even simpler example to create globally rigid (and universally rigid) configurations. (This reminds one of some of the “catastrophe” techniques in [Heaton and Timme(2022)], although Catastrophe Theory, itself, does not seem helpful.) An example that we cannot solve is also provided(Figure 20), together with other unsolved natural questions.
2 Convex Braced Polygons
We propose a class of planar frameworks, motivated from Cauchy’s Theorem in dimension three, where it may be easier to determine their infinitesimal rigidity, and even global rigidity.
Suppose that a strictly convex polygon in the Euclidean plane has edges with fixed length and some additional edges, also of fixed length, connecting some of the polygon vertices. We call this a convex braced polygon. We are concerned here when this bar framework is rigid in four senses, infinitesimally rigid, convexly rigid, globally rigid, and universally rigid. We call $(G,{\bf p})$ a bar framework, where $G$ is a finite graph whose vertices correspond to a configuration ${\bf p}=(p_{1},\dots,p_{n})$ of points $p_{i}$, $i=1\dots,n$ in some Euclidean space. The edges of $G$ correspond to fixed length bars, but not necessarily of the same length, connecting the vertices of $(G,{\bf p})$. A convex polygon in the plane is strictly convex if each vertex $p_{i}$ is in a line that intersects the polygon only at $p_{i}$, and for a convex braced polygon we assume that the vertices $(p_{1},\dots,p_{n})$ form a convex polygon with each edge $i,i+1$, modulo $n$, corresponding to a boundary edge of the polygon. When we say $(G,{\bf p})$ and $(G,{\bf q})$ have corresponding bars the same length, we mean that for each $i,j$ a bar of $G$, $\|p_{i}-p_{j}\|=\|q_{i}-q_{j}\|$, and we assume that they are both in clockwise order or both in counter-clockwise order.
Definition 2.1.
For a bar framework $(G,{\bf p})$ in $\mathbb{R}^{d}$, and ${\bf p}^{\prime}=(p^{\prime}_{1},\dots,p^{\prime}_{n})$, regarded as a configuration of vectors $p^{\prime}_{i}$ in $\mathbb{R}^{d}$, we say ${\bf p}^{\prime}$ is an infinitesimal flex of $(G,{\bf p})$, when for each edge $i,j$ of $G$
$({\bf p}_{i}-{\bf p}_{j})\cdot({\bf p}_{i}^{\prime}-{\bf p}^{\prime}_{j})=0$
We say that such a ${\bf p}^{\prime}$ is trivial if it is the derivative of a smooth family of motions of congruences of $\mathbb{R}^{d}$ restricted to ${\bf p}$. A bar framework $(G,{\bf p})$ in $\mathbb{R}^{d}$ is infinitesimally rigid if every infinitesimal flex ${\bf p}^{\prime}$ of $(G,{\bf p})$ is trivial.
Definition 2.2.
A braced strictly convex polygon $(G,{\bf p})$ in $\mathbb{R}^{2}$ is convexly rigid if every other strictly convex braced polygon $(G,{\bf q})$, whose corresponding edge lengths are the same, is congruent to $(G,{\bf p})$. (Two frameworks are congruent if $\|p_{i}-p_{j}\|=\|q_{i}-q_{j}\|$, for all $1\leq i<j\leq n$.)
Definition 2.3.
A bar framework $(G,{\bf p})$ in $\mathbb{R}^{d}$ is globally rigid in $\mathbb{R}^{d}$ if every other bar framework $(G,{\bf q})$ in $\mathbb{R}^{d}$, whose corresponding edge lengths are the same, is congruent to $(G,{\bf p})$.
Definition 2.4.
A bar framework $(G,{\bf p})$ in $\mathbb{R}^{d}$ is universally rigid if every other bar framework $(G,{\bf q})$ in $\mathbb{R}^{D}$, for all $D\geq d$, whose corresponding edge lengths are the same, is congruent to $(G,{\bf p})$.
Notice that for the graph $G$ of a braced polygon, the ordering of the vertices is understood as part of the graph $G$. With this in mind we have the following:
Theorem 2.1.
For a given braced polygon graph $G$, all strictly convex braced polygon frameworks $(G,{\bf p})$ are infinitesimally rigid, if and only if they are all convexly rigid.
One direction of this theorem follows from the following, which is related to ideas of Minkowski and extends easily to tensegrities by keeping track of the inequalities:
Proposition 2.2.
For any non-trivial infinitesimal flex ${\bf p}^{\prime}$ of a bar framework $(G,{\bf p})$, whose affine span of the vertices of ${\bf p}$ is all of $\mathbb{R}^{d}$, then $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ have corresponding bar lengths the same, but are not congruent.
Proof. If $(G,{\bf p}^{\prime})$ is an infinitesimal flex of $(G,{\bf p})$, and $i,j$ corresponds to a bar of $G$, then $(p_{i}-p_{j})\cdot(p^{\prime}_{i}-p^{\prime}_{j})=0$, and so
$\|(p_{i}+p^{\prime}_{i})-(p_{j}+p^{\prime}_{j})\|^{2}=\|(p_{i}-p_{j})+(p^{\prime}_{i}-p^{\prime}_{j})\|^{2}=$
$(p_{i}-p_{j})^{2}+(p^{\prime}_{i}-p^{\prime}_{j})^{2}+2(p_{i}-p_{j})\cdot(p^{\prime}_{i}-p^{\prime}_{j})=$
$(p_{i}-p_{j})^{2}+(p^{\prime}_{i}-p^{\prime}_{j})^{2}-2(p_{i}-p_{j})\cdot(p^{\prime}_{i}-p^{\prime}_{j})=$
$\|(p_{i}-p_{j})-(p^{\prime}_{i}-p^{\prime}_{j})\|^{2}=\|(p_{i}-p^{\prime}_{i})-(p_{j}-p^{\prime}_{j})\|^{2},$
and thus $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ have corresponding bar lengths the same. When ${\bf p}^{\prime}$ is non-trivial, and the affine span of ${\bf p}$ is all of $\mathbb{R}^{d}$, then for some pair of vertices corresponding to $i,j$, $(p_{i}-p_{j})\cdot(p^{\prime}_{i}-p^{\prime}_{j})\neq 0$, and the above calculation shows that $(G,{\bf p}+{\bf p}^{\prime})$ and $(G,{\bf p}-{\bf p}^{\prime})$ are not congruent, since the middle equality, above, is violated. ∎
One can see how this works from Figure 1. Notice that, due to the symmetry involved, the two non-congruent braced polygons on the top are mirror images of each other, but that congruence does not preserve the vertex labeling. This symmetry can be broken with different examples.
Figure 1: The two braced top polygons have corresponding bar lengths the same, but are not congruent. The bottom braced polygon, for the same graph, has an infinitesimal flex, where each non-zero $p^{\prime}_{i}$ is indicated by a small arrow. This provides the two graphs on top.
Next we need some lemmas related to convex polygons.
Lemma 2.1.
Let $f,g:\mathbb{R}^{1}\rightarrow\mathbb{R}^{1}$ be two piecewise constant, upper-semi-continuous, monotone increasing functions, with a finite number of values on $[0,2\pi)$, and $l>0$, such that for all real $x$, $f(x+l)=f(x)+2\pi$ and $g(x+l)=g(x)+2\pi$. Then there is a fixed constant $\theta$ such that $|f(x)-g(x)-\theta|<\pi$.
Proof. Notice that by replacing, say $f(x)$ by $f(x)+c_{1}$, and $g(x)$ with $g(x)+c_{2}$ where $c_{1},c_{2}$ are any real constants, it does not change the conditions on $f$ and $g$. From the period conditions, $f(x)-g(x)$ is periodic with period $l$, as well as being upper semi-continuous. From the piecewise constant conditions, there is a minimum point $m$ and a maximum point $M$ such that for all $x$ in $[m,m+l)$ and $M$ in $[m,m+l)$,
$f(m)-g(m)\leq f(x)-g(x)\leq f(M)-g(M),$ (1)
with at least one of the inequalities being strict. This is since if both inequalities are equal for all $x$, then
$f(m)-g(m)=f(x)-g(x)=f(M)-g(M)$
for all $x$, and $f(x)=g(x+Nl)$ for all $x$ for some integer $N$, and the conclusion holds easily. So, henceforward, we will assume that one of the inequalities of (1) is strict.
Then, by the inequality above, the periodicity, and the assumptions on $f$ and $g$, we have
$0<f(M)-g(M)-(f(m)-g(m))=f(M)-f(m)-(g(M)-g(m))\leq 2\pi-0=2\pi.$
Replace $f(x)$ with $F(x)=f(x)-f(m)$, and $g(x)$ with $G(x)=g(x)-g(m)$. So inequality (1) becomes
$0\leq F(x)-G(x)\leq f(M)-f(m)-(g(M)-g(m))=F(M)-G(M)<2\pi.$
In other words, the difference between the maximum value and the minimum value of $F(x)-G(x)$ is strictly less than $2\pi$. Then let
$\theta=(F(M)-G(M))/2$
the average value of the extreme values of $F(x)-G(x)$. This is what was desired, since the mid-point in an interval is no further to every point in the interval than half the length of the interval, and $\theta$ varies continuously with $f$ and $g$. ∎
Figure 2: These are two oriented pairs of line segments joined head to tail with the difference between the directions shown.
Figure 3: This shows two convex angles, i.e. oriented clockwise from $1$ to $2$, such that their average is not convex. But the $b_{2}$ bar direction has been rotated more than $\pi$, and this causes the average angle to be not convex.
Up to now, effectively we have been considering the curvature at the vertices of a polygon, as in Figure 2. Now consider the two strictly convex polygons themselves, with $n$ vertices, $(G,{\bf p})$ and $(G,{\bf q})$ with corresponding sides equal. For a point $x$ on the segment $[p_{i},p_{i+1})$, let $\theta_{i}=f(x)$, and $\phi_{i}=g(x)$ be the angle that an edge vector makes with a fixed direction in the plane so that as one proceeds around the polygons $(G,{\bf p})$ and $(G,{\bf q})$, the $i$-th angle is the direction $\theta_{i}$ of $p_{i+1}-p_{i}$ and the direction $\phi_{i}$ of $q_{i+1}-q_{i}$, respectively, both modulo $2\pi$. We make it so that both $\theta_{i}$ and $\phi_{i}$ are monotone increasing, when the orientation of the polygon is in the clockwise sense, so that $\theta_{i}<\theta_{i+1}<\theta_{i}+\pi$, $\phi_{i}<\phi_{i+1}<\phi_{i}+\pi$, and defined for all integers $i=\dots,-1,0,1,\dots$. The restriction to $\pi$ is because of the local convexity assumption on $(G,{\bf p})$ and $(G,{\bf q})$. Also for all $i$, $\theta_{i+n}=\theta_{i}+2\pi$ and $\phi_{i+n}=\phi_{i}+2\pi$, where $G$ has $n$ vertices. The maps $f$ and $g$ “unwrap” the the map from the boundary of the polygon to the circle to a similar map from the line to the line, as a function of the curve length.
The condition for strict local convexity for $(G,{\bf p})$ is that $\theta_{i}<\theta_{i+1}<\theta_{i}+\pi$, and, similarly, for $(G,{\bf q})$, it is that $\phi_{i}<\phi_{i+1}<\phi_{i}+\pi$. Even when both polygons are convex at the given corresponding vertices, Figure 3 shows that the angle for the average may not be convex for the given orientation. This is because the $b_{2}$ bar has been rotated clockwise more than $\pi$ to $c_{2}$. Local convexity can only be destroyed when one of the edges in ${\bf p}+{\bf q}$ has the opposite orientation, and this only happens when the angles between corresponding edges is $\pi$ or greater. (Note that, for computing the average configuration, it is a translate of the average of any translates of $p$ and $q$. This allows more freedom in drawing averages.) We apply Lemma 2.1 to this situation, and when $l$ is the perimeter of the polygon, and we define a piecewise constant function for both $(G,{\bf p})$ and $(G,{\bf q})$ to obtain monotone increasing piecewise constant functions.
Proof of Theorem 2.1: We have shown that when there is a non-trivial infinitesimal flex, there are non-congruent strictly convex bar frameworks $(G,{\bf p})$ and $(G,{\bf q})$. Here we assume that there are two non-congruent strictly convex braced bar polygons $(G,{\bf p})$ and $(G,{\bf q})$ with corresponding bars of the same length, both oriented clockwise, or both oriented counter-clockwise. Thus, there is piecewise-linear map from the boundary polygon of $(G,{\bf p})$ to the boundary polygon of $(G,{\bf q})$, where each boundary is of length $l$. For a point in the relative interior of each edge $p_{i},p_{i+1}$ and at $p_{i}$, define $f(x)$ to be the oriented angle in the direction $(p_{i+1}-p_{i})/||p_{i+1}-p_{i}||$ as an angle on the unit circle. Notice that $f$ is upper semi-continuous. Similarly, define $g$ for $(G,{\bf q})$. (Notice that $f$ and $g$ are a sort of piecewise Gaus map for the convex polygons.)
By Lemma 2.1 one can add an angle $\theta$ to the function $f$, which is equivalent to rotating $(G,{\bf p})$ by $\theta$ so that the average $(G,({\bf p}+{\bf q})/2)$ is strictly convex. Then ${\bf p}-{\bf q}$ is a non-trivial infinitesimal flex of $(G,({\bf p}+{\bf q})/2)$. ∎
Remark 2.3.
Even when one strictly convex braced polygon $(G,{\bf p})$ has one of its edges parallel to the corresponding edge of $(G,{\bf q})$, then the average $(G,({\bf p}+{\bf q})/2)$ may not be convex as seen in Figure 4.
Figure 4: On the top figures, this shows two strictly convex braced polygons with corresponding sides the same length, and one pair $\{1,2\}$ of corresponding sides that are parallel in the same direction, but their average is not a convex polygon. The bottom figures show the same two strictly convex braced polygons but with one rotated so that their average is convex. The dotted line segment between corresponding vertices have the average vertices at their midpoint.
Remark 2.4.
Figure 5 shows two corners where one pair $b_{1},c_{1}$ have equal length with one orientation, but the other pair $b_{2},c_{2}$ do not have equal length, but have the same orientation. However, the average pair $(b_{1}+c_{1})/2,(b_{2}+c_{2})/2$ have the opposite orientation. So the sign of the angle between the successive bars can change when corresponding bars do not have the same length.
Figure 5: These are two corners, where not all corresponding sides are equal, and the average configuration has the opposite orientation
Remark 2.5.
Figure 6 shows two bar frameworks that have corresponding bars of the same length. The one on the right is a strictly convex braced polygons that is infinitesimally rigid in all strictly convex realizations and so it is convexly rigid by Theorem 2.1. On the other hand, even though the graph of these polygons are globally rigid in the plane for all generic configurations, due to the configuration on the left, these configurations are necessarily not globally rigid in the plane and the one on the left is necessarily not strictly convex.
Figure 6: This shows two non-congruent bar frameworks for the same graph that have corresponding bars the same length. The one on the right is a braced strictly convex polygon. The one in the middle is an average of the two other non-congruent configurations. Note that the middle framework is convex, but not strictly convex, and is necessarily not infinitesimally rigid in the plane. The points $1,7,6,5$ are fixed in all three realizations.
3 Braced Polygons with Proper Stresses
In this section, we shall strengthen Theorem 2.1 by restricting our attention to special families of braced polygons. Our discussion is motivated from the result by Geleji and Jordan
[Geleji and Jordán(2013)], which characterizes the global rigidity of all strictly convex realizations when the underlying graph is a circuit in the generic rigidity matroid. We shall develop a novel analysis of the rigidity of braced polygons by focusing on minimal 3-connectivity.
3.1 Proper Stresses and Super Stability
Our tool to prove global rigidity (or even universal rigidity) is super stability based on self-stresses.
Definition 3.1.
For a configuration $(G,{\bf p})$, a stress is a function that maps the set of edges to real numbers, $\omega:E\to\mathbb{R}$. The stress is in equilibrium if $\sum_{j}\omega_{ij}(\mathbf{p}_{i}-\mathbf{p}_{j})=\mathbf{0}$ for every vertex $i$. A braced polygon is properly stressed if it has an equilibrium stress that is positive on the boundary polygon and negative on the interior edges. The stress matrix $\Omega$ of a stress $\omega$ is an $n\times n$ symmetric matrix defined as $\Omega_{ij}=-\omega_{ji}$ for each edge $e_{ij}$, $\Omega_{ij}=0$ if $i\not=j$ and there is no edge between $v_{i}$ and $v_{j}$, and that each row sums to 00.
Definition 3.2.
A framework $(G,{\bf p})$ of dimension $d$ with stress matrix $\Omega$ is super stable if $\Omega$ is positive semidefinite with rank $n-d-1$, and all affine transformations that preserve the length of all edges are trivial motions of the framework.
A basic theorem in [Connelly and Guest(2022), Theorem 5.14.1] states that super stable frameworks are universally rigid. The reverse is not true. See Figure 21E in the plane, and see [Connelly and Gortler(2015)] for the general situation.
For braced polygons, the following sufficient condition for super stability is useful.
Theorem 3.1 (Connelly [Connelly(1982)]).
Suppose a strictly convex braced polygon $(G,{\mathbf{p}})$ is properly stressed. Then it is super stable for the corresponding stress matrix.
The following easy observation that follows from convexity is another key tool.
Proposition 3.2.
Let $(G,{\mathbf{p}})$ be a strictly convex braced polygon and suppose that $(G,{\mathbf{p}})$ has a non-zero equilibrium stress $\omega$ that is non-positive on all the interior edges. Then $\omega$ is positive on the boundary edges.
Proof. At each vertex consider the three vectors, centered at the vertex; the sum of the interior edges weighted by their non-positive stresses, and the two boundary edges weighted by their stresses. The equilibrium condition implies that the sum of these three vectors is 00. If just one of these vectors is 00, it implies that the other two are negatives of each other, and this contradicts the strict convexity at that vertex. Similarly, if two of the vectors are 00, so is the third and they are all 00. Then the local strict convexity at the vertex implies that when all three are non-zero, the boundary vertex stresses are both positive by the independence of the three vectors.
So either all the stresses are 00, or all the boundary stresses are positive, and there is at least one interior stress that is negative at each vertex. ∎
3.2 Minimally 3-connected Braced Polygons with Proper Stresses
In view of Theorem 3.1, the next natural question would be to identify strictly convex braced polygons having proper stresses. We have a concrete answer for the family of minimally 3-connected braced polygons, whose formal definition is given as follows.
Definition 3.3.
A braced polygon is minimally $k$-connected if it is $k$-connected, and removing any brace (interior edge) would cause it to be no longer $k$ vertex connected.
The following theorem is a main technical observation.
Theorem 3.3.
Every 3-connected braced polygon $P$ has a strictly convex realization that is properly stressed. In addition, if $B$ is a set of braces such that $P-B$ is 3-connected, then there exists a strictly convex realization such that $P-B$ is properly stressed and all braces in $B$ have positive stresses.
Proving Theorem 3.3 requires several additional ingredients that will be proved individually. First, we show there is an inductive procedure to produce any 3-connected braced polygon. Next, we show that if a convex but not strictly convex braced polygon is properly stressed, then there exists a nearby strictly convex configuration with desired stress by minimizing a certain energy function.
Proposition 3.4.
Every 3-connected braced polygon $P$ with at least 4 vertices can be constructed by adding one brace at a time by the following operations starting from $Q=K_{4}$:
  • remove a boundary edge $(i,j)$, add a new vertex $k$ and create two new boundary edges $(i,k)$ and $(j,k)$, and connect $k$ to another vertex.
  • remove two boundary edges (possibly with one overlapping vertex) $(i_{1},j_{1})$ and $(i_{2},j_{2})$, add four boundary edges $(i_{1},k_{1}),(j_{1},k_{1}),(i_{2},k_{2}),(j_{2},k_{2})$ and a brace $(k_{1},k_{2})$.
  • add a brace between two existing vertices.
The graph $Q$ is 3-connected in every step. Furthermore, if $P$ is minimally 3 vertex connected, then it is constructed by only using the first 2 operations.
Proof. Since $P$ is 3-connected, every brace must be crossed by another brace in the interior of the polygon and all braces must be connected by crossing points. Take 2 braces that cross each other and identify them with the 2 braces in $Q=K_{4}$ together with 4 vertices.
Figure 7: The degree 2 vertices are shown only to make the vertex correspondence clearer. They don’t exist in our construction until they are added through a new brace. In each step, one brace that crosses existing braces is added, so the new graph must stay 3 vertex connected.
If $P\not=Q$, then there exists a brace $(k_{1},k_{2})$ from $P$ that crosses a brace identified in $Q$. Otherwise, the braces in $P$ identified in $Q$ and the remaining braces never cross each other, contradicting $P$ being 3-connected. Depending on the number of vertices in $k_{1}$ and $k_{2}$ identified in $Q$, choose one of the 3 operations so that the new edge $(k_{1}^{\prime},k_{2}^{\prime})$ has the same crossing pattern with the existing braces in $Q$ as $(k_{1},k_{2})$ with the corresponding braces in $P$. This procedure is demonstrated in Figure 7.
To prove the last statement, we show that if we do the last operation, then the added edge will never become necessary for the 3-connectivity. Therefore, $P$ is not minimal if the last operation is used in the process.
Suppose that we can add $(k_{1},k_{2})$ to $Q$ by the last operation, then the new graph is immediately no longer minimal. Since $Q$ is 3-connected without $(k_{1},k_{2})$, there must be a path $L$ from $k_{1}$ to $k_{2}$ in the interior of the polygon joined by the crossings without $(k_{1},k_{2})$. Therefore, any edge added later that crosses $(k_{1},k_{2})$ must also cross the path $L$. As a result, all the braces are still connected by crossings, even if we remove $(k_{1},k_{2})$. ∎
Proposition 3.5.
There exists a differentiable non-negative strictly monotonically increasing real function $f$ defined on $[0,\infty)$ such that given $a,\omega,m,\delta>0$, $f(a)=\frac{1}{m}$, $f^{\prime}(a)=\omega$, and $f(a+\delta)>1$. Similarly, there exists a differentiable positive strictly monotonically decreasing function $g$ defined on $(0,\infty)$ such that given $a,m,\delta>0$ and $\omega<0$, $g(a)=\frac{1}{m}$, $g^{\prime}(a)=\omega$, and $g(a-\delta)>1$.
[…]
Figure 8: This is the derivative of $f$. The red region has area $\frac{1}{m}$. The blue region has area at least $1$. The rest of the function only has to be positive and continuous.
[…]
Figure 9: brace $(5,6)$ is added by the second operation defined in Proposition 3.4
Now we are ready to prove Theorem 3.3.
Proof of Theorem 3.3: We proceed by induction. Since $P$ is minimally 3-connected, it can be constructed using 2 operations according to Proposition 3.4. The base case, $K_{4}$, can be placed on a square to generate a stress with the desired signs. For the induction step, we need to show that if a braced polygon $P$ is properly stressed in a strictly convex configuration, then the braced polygon $Q$ after adding a brace $(k_{1},k_{2})$ through an operation can also be properly stressed in a strictly convex configuration. Figure 9 is a case where the brace $(5,6)$ is added.
The same embedding as $P$ with new vertices placed on existing edges (shown in Figure 9) immediately gives an equilibrium stress $\omega$ on $Q$ where all forces (stress times length) on existing edges remain unchanged and the stress on $(k_{1},k_{2})$ is 00. We call this configuration $(Q,\mathbf{p}_{0})$. There are two problems that need to be solved: $(Q,\mathbf{p}_{0})$ is no longer strictly convex and the stress is not strictly negative on the new brace.
Let $p_{i}$ be the coordinate of the vertex $i$ in $(Q,\mathbf{p})$, $l(\mathbf{p})$ be the length function maps a configuration $\mathbf{p}$ to $\mathbb{R}^{m}$, $l_{ij}(\mathbf{p})$ be the distance from $p_{i}$ to $p_{j}$, $m$ be the number of edges in $Q$, $\epsilon,\delta>0$ be real numbers, $\omega_{ij}$ be the stress on the edge $(i,j)$ with the only exception $\omega_{k_{1}k_{2}}=-\epsilon$. By Proposition 3.5, we can find monotonic and differentiable energy functions $E_{ij}(l^{2}_{ij})$ for each edge $(i,j)$ such that $E_{ij}=\frac{1}{m}$ at $\mathbf{p}_{0}$, $E_{ij}^{\prime}(l_{ij}^{2})=\omega_{ij}$, and $E_{ij}>1$ if $l_{ij}^{2}$ increases by $\delta$ for $\omega_{ij}>0$ or decreases by $\delta$ for $\omega_{ij}<0$. Let $E(Q,\mathbf{p})=\sum_{ij\in Q}E_{ij}(l_{ij}^{2}(\mathbf{p}))$. This energy function has a few good properties:
  • At the given configuration $(Q,\mathbf{p}_{0})$, $E(Q,\mathbf{p}_{0})=\sum_{ij\in Q}\frac{1}{m}=1$
  • Fixing a vertex $i$, $\frac{\partial}{\partial x_{i}}E_{ij}(l_{ij}^{2})=2E_{ij}^{\prime}(l^{2}_{ij})(x_{i}-x_{j})$, hence a critical point of $E$ gives an equilibrium stress with value $E_{ij}^{\prime}(l^{2}_{ij})$ on edge $(i,j)$
  • Pulling $k_{1}$ and $k_{2}$ outwards infinitesimally in direction orthogonal to the boundary will decrease $E$, hence there exists a configurations nearby with $E<1$.
Now we make the following observation: if $\delta$ is sufficiently small, then $l_{ij}(\mathbf{p})$ must stay close to $l_{ij}(\mathbf{p}_{0})$ in order to keep $E(Q,\mathbf{p})<1$. Suppose that a boundary edge $(i,j)$ decreases in length by $\xi$. Because the configuration $(Q,\mathbf{p}_{0})$ is super stable, $\xi$ must converge to 00 as $\delta$ goes to 00. Therefore, we find that $l_{ij}(\mathbf{p})\to l_{ij}(\mathbf{p}_{0})$ for all $\mathbf{p}$ such that $E(Q,\mathbf{p})<1$ as $\delta\to 0$. The argument for braces is similar. This implies that if $\delta$ is small enough, there is a critical point $l(\mathbf{q})$ of $E$ in the small neighborhood of $l(\mathbf{p}_{0})$ giving the minimum of $E$.
Next, we claim that if $l(\mathbf{q})$ is close to $l(\mathbf{p}_{0})$, then the configuration $\mathbf{q}$ must be close to $\mathbf{p}_{0}$ (up to Euclidean isometry). Consider the sequence $\delta_{i}=\frac{1}{i}$. Without loss of generality, we fix a vertex and the direction of an edge connected to it. Let $\{\mathbf{q}_{i}\}=\mathbf{q}_{1},\mathbf{q}_{2},...$ be a sequence of possible configurations. Since the lengths are all bounded, a configuration will always stay in a compact neighborhood. Therefore, some subsequence of $\{\mathbf{q}_{i}\}$ converges to a point $\tilde{\mathbf{p}}$. If $\tilde{\mathbf{p}}\not=\mathbf{p}_{0}$, then we have another configuration with $l(\tilde{\mathbf{p}})=l(\mathbf{p}_{0})$, which contradicts super stability.
Finally, for a chosen $\epsilon$, pick a small enough $\delta$, and let a critical point of $E$ near $\mathbf{p}_{0}$ be $\mathbf{q}$. $E^{\prime}(l^{2}(\mathbf{q}))$ is an equilibrium stress. If $\delta$ is sufficiently small, no stress from $P$ can change sign because $E^{\prime}$ is continuous. Notice that the stress on $(k_{1},k_{2})$ must be negative because our $E_{k_{1}k_{2}}$ strictly decreases in Proposition 3.5.
To see that $\mathbf{q}$ is strictly convex, all vertices are in a small neighborhood of $\mathbf{p}$, therefore, every vertex that was strictly convex stays so with a sufficiently small $\delta$. For the newly added vertices, they have degree 3 with positive stress on the boundary and negative stress on a brace, so they must be strictly convex. This concludes the last part of Theorem 3.3.
By the same energy minimization argument, if an edge is not necessary for the 3 vertex connectivity, then it can have either positive or negative stress. We can pick $\epsilon$ to be either positive or negative in this case and a small $\delta$. A point in the small neighborhood of a strictly convex polygon must also be strictly convex. ∎
3.3 Minimal 3-connectivity and circuits
Minimally 3-connected braced polygons also enjoy the following special property.
Lemma 3.1.
Let $(G,{\mathbf{p}})$ be a minimally 3-connected strictly convex polygon and $\omega$ be a non-zero equilibrium stress of $(G,{\mathbf{p}})$ that is non-positive on the interior braces. Then $\omega$ is a proper stress and the space of equilibrium stresses of $(G,{\mathbf{p}})$ has dimension one.
Proof. By Proposition 3.2, $\omega$ is positive on the boundary edges. Suppose $\omega$ is zero on some interior brace. Let $Z$ be the set of interior braces which have zero stress in $\omega$. Then, $\omega$ is a proper stress of $(G-Z,{\mathbf{p}})$. This implies from Theorem 3.1 that $G-Z$ is 3-connected. Since $Z$ contains at least one brace, this contradicts the minimal 3-connectivity of $G$. Thus $\omega$ is a proper stress.
Suppose that the space of equilibrium stresses of $(G,{\mathbf{p}})$ has dimension at least two. Let $\omega_{1}$ be a proper stress and $\omega_{2}$ be an arbitrary equilibrium stress that is not a scalar multiple of $\omega_{1}$. Consider $\omega_{t}:=\omega_{1}+t\omega_{2}$ for a parameter $t$ starting at $t=0$, and increase $t$ continuously. Let $\bar{t}$ be the smallest $t$ such that $\omega_{t}(e)=0$ holds for some edge $e$. Then the existence of $\omega_{\bar{t}}$ contradicts the former part of the statement. ∎
Generic rigidity circuits are one of the central classes of graphs in the context of global rigidity.
Definition 3.4.
A graph $G$, with with $n$ vertices and $m$ edges, is called a (planar) circuit if $m=2n-2$ and for every subgraph on $k$ vertices such that $2\leq k<n$, the number of edges is smaller than or equal to $32k-$ 3.
An important corollary of Theorem 3.3 and Lemma 3.1 is the following.
Lemma 3.2.
Let $G$ be a minimally 3-connected polygon graph with $n$ vertices and $m$ edges. If $m\geq 2n-2$, then $G$ is a circuit. Moreover, every strictly convex framework $(G,{\mathbf{p}})$ is properly stressed, and $m=n-2$.
Proof. By Theorem 3.3, there is a strictly convex framework $(G,{\mathbf{q}})$ having a proper stress. By Lemma 3.1, the space of equilibrium stresses of $(G,{\mathbf{q}})$ has dimension one. Therefore, by $m\geq 2n-2$, we must have $m=2n-2$ and the rigidity matrix has rank equal to $32n-$ 3. Moreover, every edge has a non-zero stress in the proper stress, which means that $G$ is a circuit.
We now show that every strictly convex polygon $(G,{\mathbf{p}})$ is properly stressed. Let $G_{1},\dots,G_{s}$ be the set of generically minimally rigid spanning subgraphs of $G$, and let $N_{i}$ be an open subspace of the configuration space of strictly convex realizations of $G$ such that $G_{i}$ forms a minimally infinitesimally rigid framework $(G_{i},{\mathbf{p}})$ for any ${\mathbf{p}}\in N_{i}$. Over $N_{i}$, the rigidity matrix can be solved with respect to the row vectors associated with $G_{i}$ (after an appropriate pin down), so the space of equilibrium stresses can be described as a rational function in the entries of ${\mathbf{p}}$.
As we have seen above, $G$ has an infinitesimally rigid strictly convex realization $(G,{\mathbf{p}}_{0})$ having a proper stress $\omega_{0}$. Consider any strictly convex framework $(G,{\mathbf{p}}_{1})$ of $G$. Since the configuration space of a convex polygon with fixed graph is connected, we can continuously deform $(G,{\mathbf{p}}_{0})$ into $(G,{\mathbf{p}}_{1})$ by a continuous path ${\mathbf{p}}_{t}\ (t\in[0,1])$ within the space of convex polygons. Suppose that $(G,{\mathbf{p}}_{t})$ is infinitesimally rigid for any $t\in[0,1]$, i.e., the continuous path is covered by $\bigcup_{i=1}^{s}N_{i}$. Then there is a non-zero equilibrium stress $\omega_{t}$ of $(G,{\mathbf{p}}_{t})$ such that the stress varies continuously in $t$. Then, by Lemma 3.1 and the fact that $\omega_{0}$ is a proper stress, $\omega_{t}$ remains a proper stress for all $t\in[0,1]$. In particular, $(G,{\mathbf{p}}_{1})$ is properly stressed.
Suppose $(G,{\mathbf{p}}_{t})$ is not infinitesimally rigid. Since $(G,{\mathbf{p}}_{0})$ is infinitesimally rigid, we have $t>0$. We may assume that $(G,{\mathbf{p}}_{t^{\prime}})$ is infinitesimally rigid and for any $t^{\prime}\in[0,t)$ and is properly stressed (by the same argument as above). We consider a sequence $\{{\mathbf{p}}_{t_{i}}\}_{i=1,2.\dots}$ of point configurations that converges to ${\mathbf{p}}_{t}$. Let $\omega_{t_{i}}$ be a proper stress of $(G,{\mathbf{p}}_{t_{i}})$ having unit norm. By the compactness of the unit sphere, a subsequence of $\{\omega_{t_{i}}\}_{i=1,2,\dots}$ converges to a nonzero $\omega$, which is an equilibrium stress of $(G,{\mathbf{p}}_{t})$. Since $\omega$ is non-positive on the interior braces, Lemma 3.1 implies that $\omega$ is a proper stress. However, since $G$ is a circuit and $(G,{\mathbf{p}}_{t})$ is not infinitesimally rigid, the space of equilibrium stressed of $(G,{\mathbf{p}}_{t})$ is at least two, contradicting Lemma 3.1. Thus also $m=n-2$ because $G$ is a circuit. This completes the proof. ∎
3.4 Main theorems
We are now ready to give our main result.
Theorem 3.6.
The following are equivalent for a minimally 3-connected braced polygon graph $G$.
  • (a)
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are infinitesimally rigid.
  • (b)
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are convexly rigid.
  • (c)
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are globally rigid.
  • (d)
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are super stable.
  • (e)
    All strictly convex polygon frameworks $(G,{\mathbf{p}})$ are properly stressed.
  • (g)
    $G$ is a circuit.
  • (h)
    $G$ has $n$ boundary vertices and $n-2$ internal braces ($n\geq 4$).
Proof. The equivalence between (a) and (b) follows from Theorem 2.1. (c) implies (b) by definition, and (d) implies (c) by Connelly’s super stability theorem. (e) implies (d) by Theorem 3.1. (g) implies (e) by Lemma 3.2. Finally, Suppose (a) holds. By Theorem 3.3, $G$ has an infinitesimally rigid strictly convex framework $(G,{\mathbf{p}})$ with a proper stress. Hence, $m\geq 2n-3+1=2n-2$. By Lemma 3.2, $G$ is a circuit and (g) follows. Similarly, Lemma 3.2 shows that (g) and (h) are equivalent. ∎
Geleji and Jordán [Geleji and Jordán(2013), Theorem 1.2] gave a criterion, known as the “unique interval property”(See Appendix), that is equivalent to all convex configurations of a polygonal circuit having proper stress, we show that minimal 3 connectedness is necessary and sufficient.
Theorem 3.7.
Let $G$ be a braced polygonal graph with $n$ vertices and $22n-$ 2 edges. Then, $G$ is minimally 3-connected if and only if all strictly convex polygon frameworks $(G,{\mathbf{p}})$ are properly stressed.
Proof. Suppose $G$ is minimally 3-connected. By Lemma 3.2, $G$ is a circuit. By Theorem 3.6, all strictly convex polygon frameworks of $G$ are properly stressed.
To see the other direction, suppose that $G$ is not minimally 3-connected. Then, pick a spanning minimally 3-connected subgraph $H$ of $G$. By Theorem 3.3, there is a strictly convex framework $(H,{\mathbf{p}})$ which is properly stressed. If $(G,{\mathbf{p}})$ is not infinitesimally rigid, then we are done by Theorem 2.1. So, assume $(G,{\mathbf{p}})$ is infinitesimally rigid. Then the rank of the rigidity matrix of $(G,{\mathbf{p}})$ is $32n-$ 3 and the space of equilibrium stresses of $(G,{\mathbf{p}})$ is one dimensional by $m=2n-2$. This in turn implies that no edge in $E(G)\setminus E(H)$ can be stressed in $(G,{\mathbf{p}})$ (i.e., each edge in $E(G)\setminus E(H)$ is a coloop in the rigidity matroid of $(G,{\mathbf{p}})$). In other words, $(G,{\mathbf{p}})$ is not properly stressed. ∎
Remark 3.1.
The condition of having $n-2$ braces guarantees the existence of a stress in any configuration. Then the minimal 3-connectivity prevents the stress on any edge from changing its sign. Does having a stress in a minimally 3 connected convex braced polygon imply it is properly stressed in that particular configuration when there are not enough braces? The answer is negative, as shown in Figure 10. This shows that it is necessary for the stress to vary continuously with respect to the configuration.
Figure 10: This braced polygon is minimally 3 connected. However, $e_{47}$ has zero stress when vertices $\{1,2,3,4,5,6\}$ are placed on a circle and vertices $\{6,7,8,9,10,1\}$ are placed on a circle. The set of convex stressed configurations is disconnected for the braced polygon graph so we cannot vary the stress continuously from a properly stressed configuration to this figure.
In order to interpolate between our result and the work of Geleji and Jordán, in the appendix we give a combinatorial proof for the equivalence between the unique interval property and minimal 3-connectivity for circuits.
Figure 11: Relation among properties of braced polygonal graphs.
You might wonder if minimal 3-connectivity is equivalent to other properties listed in Theorem 3.6 for circuits. There are non-minimally 3-connected circuits which satisfy (a), (b), (c), or (d) in Theorem 3.6. Examples are given in the next subsection. Figure 11 summarizes the results.
3.5 Braced Polygon Graph Circuits with 7 or 8 vertices
As a concrete example and application of our theorems, in this subsection we shall examine strictly convex braced polygons with $7$ vertices and $12$ edges. Figures 1214, and 15 show the examples of strictly convex braced polygons, with $7$ vertices and $12$ edges, collected into the catagories of being always super stable with a proper stress (Figure 12), always super stable, but with some configurations that have cables for internal braces (Figure 14), and those that are not globally rigid in the plane, but are convexly rigid in the plane (Figure 15).
Figure 12: These are all the $7$-convex braced circuits with a proper stress. All the internal braces are struts in all configurations. See Theorem 3.7.
Figure 13: These are all the $8$-convex braced circuits with a proper stress. All the internal braces are struts in all configurations. See Theorem 3.7.
Figure 14: These are all the graphs of braced 7-gons, where the configuration is super stable in all strictly convex realizations, but where some of the (braced) edges of the configurations are not always struts. Each red edge may be a cable instead of strut in some configurations.
Note that the red edges in Figure 14 are defined as those internal bracing edges that do not always have a negative sign for strictly convex configurations. It is necessary, in that case, that the edge does not have a degree $3$ vertex, at one end, but it is not always sufficient. Indeed, in the last two cases in the second column, there is an edge that has both vertices of degree $3$, but it is not colored red. In order for a stress on an edge to remain negative for all convex configurations, from Theorem 3.7, it is necessary and sufficient, that, when the edge is removed, the graph should fail to be $3$-connected.
Figure 15: These are all the $7$-gon braced circuits in the plane, which are convexly rigid, i.e. unique in all convex realizations, but are not globally rigid in the plane. The blue vertices may have realizations in the plane, where they are inside the convex hull of other vertices and, in that case, the configuration is not globally rigid in the plane. The idea is to move the green vertices to the interior of the corresponding red edge, and then there is a $2$-dimensional space of stresses, indicating a non-trivial infinitesimal motion, such that one of the implied equivalent but not congruent configurations is convex and the other is not. See Figure 6, which is related to the second framework here.
4 Examples and Questions
For a specific configuration, global rigidity is known to be NP-hard, while infinitesimal rigidity can be determined by computing the null space of a matrix. According to Theorem 2.1, these two are identical in the set of strictly convex configurations for braced polygon graphs. A question is whether this fact can be used to design an efficient algorithm.
Question 4.1.
For a graph G, with a Hamiltonian cycle, that is generically rigid in the plane, is there a reasonable polynomial-time algorithm to determine when all strictly convex configurations, for the Hamiltonian cycle, are infinitesimally rigid in the plane?
Figure 16: This shows a strictly convex braced polygon that is infinitesimally rigid in all strictly convex configurations, but it is not obtained just from Henneberg Type I operations.
For example, one way to construct an example of a convex braced polygon that is infinitesimally rigid in all the strictly convex configurations is to build from a triangle inductively by connecting a new vertex of degree $2$ to the previous vertices, a Type I Henneberg operation. However, this procedure does not generate all such examples, as demonstrated in Figure 16.
Question 4.2.
Is there an analogue of Theorem 2.1 for $\mathbb{R}^{3}$ (or in higher dimensions)? Namely if all the vertices of two bar frameworks in $\mathbb{R}^{3}$, each of whose vertices lie on their convex hull, when is there a rotation of one of them such that their average is convex.
When the framework is the one skeleton of a strictly convex triangulated sphere in $\mathbb{R}^{3}$ it is known [Connelly and Guest(2022)] that framework is convexly rigid by Cauchy’s Theorem and infinitesimally rigid by Dehn’s Theorem. On the other hand, there are examples of non-convex triangulated surfaces, whose vertices lie on their convex hull (weakly convex) that are always infinitesimally rigid. See [Izmestiev and Schlenker(2010), Connelly and Schlenker(2010)]. Are they convexly rigid?
Question 4.3.
How does one tell if a braced polygon graph is universally rigid in all convex configurations?
One idea is that two graphs known to be super stable in all convex configurations can be “stuck” together through superposition to create another super stable graph. When two graphs share 3 vertices not on a line, an edge that has a positive stress can cancel with an edge with a negative stress. This is shown in Figure 17 where a heptagon is assembled by adding three smaller polygons while canceling two edges stresses. Note that here one (but not both) of the red edges could be a cable depending on the configuration.
Figure 17: An example of assembling larger super stable graphs from smaller graphs. The cancelled edges have opposite signs in all convex configurations in the components.
Sometimes, it is possible to have super stable circuit framework configurations as in Figure 18, where the vertices are in a strictly convex position, but not all the boundary edges of the corresponding convex polygon are part of the framework. Can you see how the superposition works here? In such a convex position, the cables and struts always have the corresponding positive and negative stresses, when in a strictly convex position.
Figure 18: These tensegrities are always super stable, via superposition, when their vertices are in a strictly convex configuration. The top part of Figure 19 shows the derivation for one of the cases.
Figure 19: These tensegrities are always super stable, via superposition, when their vertices are in a strictly convex configuration. Here the red edges are the ones that always cancel. Notice that the tensegrity of the bottom configuration is the same as the Grünbaum tensegrity of Figure 20, except that the $84$ and $51$ edges have been switched in the plane projection.
For all strictly convex braced polygons with seven or fewer vertices, we can decide whether there is an example that is not globally rigid in the plane as a bar framework, or in all cases it is super stable. However, we have found some strictly convex cases with $8$ vertices, where we know that the framework is not super stable, and by Theorem $12.1$ in [Connelly and Gortler(2015)] not even universally rigid in some strictly convex cases. But we cannot decide whether it is globally rigid or not for the non-universally rigid strictly convex cases in the plane. The graph of Figure 20 is such a case in the plane projection, when the outside polygon is strictly convex. It looks like the examples of Figure 20 are always globally rigid as bar frameworks in the plane, but we have no proof.
In the case of the Grünbaum bar framework (Figure 20), the $1458$ and $2367$ vertices each form an infinitesimally rigid framework in the plane, since they are constructed of two triangles with a common edge in the plane. If there is another configuration of these (infinitesimally) rigid subsets of the configuration, then the other non edges $15$ and $37$ are implicitly cables, since they can only decrease in length. If either one of these lengths does not decrease in length, then the bottom row of Figure 19 shows that with this added bar, the configuration would be universally rigid. Nevertheless, with this added information, we still don’t see how the Grünbaum configurations are always globally rigid in the plane.
Figure 20: This shows a $3$-dimensional tensegrity that projects orthogonally into the plane. It is a superposition of four planar quadrilaterals shown on the right. In three-space it is super stable, but in the plane it is only universally rigid.
Figure 21 shows that there is a region of the space of convex configurations of the Grünbaum bar framework that are not universally rigid, and other regions that are super stable, assuming a rotational symmetry of 90 degrees. The red region corresponds to the positon of vertex $7$, where the configuration is not universally rigid, but the polygon is still strictly convex. It turns out that when the point $7$ is below the $2,6$ line in Figure 21, the configuration is not universally rigid.
Figure 21: These tensegrities are various configurations of the Grübaum tensegrity in the plane as in Figure 20. They are all infinitesimally rigid as bar frameworks, but those members that have 00 stress are omitted to further understand whether the stress matrix is PSD or not. But the 00 stressed bars still contribute to the bar framework’s infinitesimal rigidity. Thus, the stresses of the stress matrix, under suitable normalization, will change continuously and so will the eigenvalues of the stress matrix.
A This essentially is a square with two struts as diagonals, and so it is super stable.
B This continues to be super stable, but $\omega_{62}<0$
C Point $7$ and all the other points lie on the same circle and so there is a
C reflection symmetry in the whole configuration that has proper stress with
C $\omega_{84}=\omega_{62}=0$ and all other stresses on the internal edges negative.
D These configurations must be super stable until Point $7$ crosses the horizontal line,
D and $\omega_{84}=\omega_{62}>0$ by looking ahead to configurations E.
E This is the Grünbaum tensegrity in Figure 20,
E so the stress matrix $\Omega$ is PSD with only one extra 00 eigen value.
F The circular arc of the red region is part of a circle with center at the midpoint
F between the points $6$ and $8$. So the angle at Point $7$ is at most a right angle and the
F configuration is convex. Looking ahead to the configuration I, the 00 eigen value became negative.
G None of the eigenvalues or stress signs change, but the region is not strictly convex
H Now the region is not convex, and none of the eigenvalues or stress signs have changed
I $\omega_{67}=0$ and so the value of the stress energy is the negative of the stress
I energy of a square with diagonals as struts.
Cases from Figure 21
Remark 4.1.
Another simple way to create examples of universally rigid frameworks is to adjoin a sequence of universally rigid “plates” together with a strut from the first to last as in Figure 22.
Figure 22: The tensegrities on the left, thought of as bar frameworks, are made of rigid triangles placed edge-to-edge with a strut connecting the first to last going through all the interior of the triangles, and thus easily seen to be universally rigid, just by the triangle inequality. This shows how to construct such a framework as a superposition of simpler tensegrities. The red and green members, when superimposed, have their stresses cancel.
One of the problems with applying this idea more generally is that each of the plates needs to be universally rigid, at least as a bar framework. If there is an open set of configurations that is universally rigid, except for a triangle in the plane, if more than one of the plates is not a triangle, it will have a non-zero stress, and then the union will have a $2$-dimensional set of stresses. We are treating the case of circuits, here, which have only a one-dimensional space of stresses. The last case here in Figure 22 is the next to last example in the first column, in the table of Figure 14.
Question 4.4.
Is there a convex configuration of Figure 20 that is not globally rigid?
In this case, it is not guaranteed, but natural, to look for nearby non-universally rigid configurations in the plane. The given configuration is universally rigid, but not super stable because it can be lifted to dimension 3. We expect some small perturbation can turn the extra zero eigenvalue negative. Indeed, it turns out that there are small perturbations in the plane, where the stress matrix picks up a negative eigenvalue, and so it is not universally rigid. A random search among strictly convex $8$-gons for this graph in the plane is unlikely to find these non-universally rigid configurations, even though they form an open subset of such configurations.
To see that the given configuration of Figure 20 is universally rigid, blue vertices are in one horizontal plane, and the red vertices are in a parallel horizontal plane. All the red struts have a stress that is the negative of the stress of the overlapping red cables, because of the symmetry. For example, the $7,8$ line is parallel to the $6,8$ line. Thus, the stresses cancel. The projection is a strictly convex octagon, and thus is not always universally rigid, from the calculation mentioned above. The projection in the plane is always infinitesimally rigid and thus the stress matrix, (and its eigenvalues) vary continuously as the configuration varies continuously in the plane. It is not clear to us whether this tensegrity is always globally rigid in the plane when it is strictly convex. This example is from Grünbaum and Shephard’s “Lectures in Lost Mathematics” Figure 18. [Grünbaum and Shephard(1975)] See also [Connelly and Guest(2022)] Figure 11.10.
The central question, possibly wishful thinking, is whether there is a way to determine global rigidity for convex braced polygonal circuits, by the methods presented here. We have a way to tell when such a circuit is NOT globally rigid by Theorem 2.1, such as the example in Figure 6. Minimally 3-connected cases can be solved using Theorem 3.7, such as those in Figure 12. When neither of the theorems apply, our best hope is to use the superposition technique in Figure 17. This technique works for most of the cases that we have tried. There are still open cases such as Figure 20. This leads to the following:
Question 4.5.
If a braced strictly convex polygonal circuit $(G,{\bf p})$ is not globally rigid in the plane, is it true that there is a, convex but possibly not strictly convex circuit, $(G,{\bf q})$, with a non-trivial infinitesimal flex ${\bf q}^{\prime}$ such that ${\bf p}={\bf q}+{\bf q}^{\prime}$?
One should keep in mind that when there is a configuration ${\bf q}$ that is convex but not strictly convex with an infinitesimal flex ${\bf q}^{\prime}$ such that both ${\bf q}+{\bf q}^{\prime}$ and ${\bf q}-{\bf q}^{\prime}$ are not strictly convex, such as in the case of Figure 22. However, if Question 4.5 is true the octagon example of Grünbaum, Figure 20 projected into the plane would always be globally rigid.
Question 4.6.
From Figure 14 with $29$ super stable and strictly convex braced polygon circuits with $7$ vertices, versus Table 12 with only $5$ graphs that are always properly stressed, it seems that there are quite a bit more examples when we allow some of the stresses to change sign. Does this persist with larger numbers of vertices? What is the asymptotic behavior as the number of vertices increases?
If the number of vertices $n$ is large, it is likely that the situation in Figure 1 will be the dominant case. There are more ways for subgraphs to be stressed in a convex shape with large $n$. If two subgraphs can be stressed at the same time, then the configuration will have two stresses. As a result, the configuration will not even be convexly rigid. If we only look at convexly rigid braced polygon circuits with $n$ vertices, then is it possible to estimate how many of them are globally rigid, universally rigid, or properly stressed in all convex configurations?
Question 4.7.
Finally, rigidity cannot be done without asking the most obvious question: is there a braced polygon that is generically rigid but flexible?
Without convexity, it is possible to place all edges of Figure 20 on a rectangular grid for flexibility. With convexity, it is not obvious. Our intuition says this should not be possible just like Cauchy’s Rigidity Theorem. Theorem 2.1 shows that a generically rigid braced polygon is closely related to convex polytopes. Grasegger and Legersky’s idea on NAC coloring can assert the existence of flexible configurations[Grasegger et al.(2017)Grasegger, Legerský, and Schicho], yet convexity rules out the affine motions they use to construct flexible configurations.
Appendix A Appendix: unique interval property
In this appendix, we shall discuss a relation between minimal 3-connectivity and the unique interval property introduced by Geleji and Jordán [Geleji and Jordán(2013)].
Without loss of generality, we assume that the vertices are labeled counterclockwise on the boundary of the outside Hamiltonian circuit for braced polygons.
Definition A.1.
We say that a braced polygon $G$ satisfies the unique interval property
[Geleji and Jordán(2013)] if
  1. 1.
    there exists a connected segment of the outside polygon $I$ of vertices $\{v_{1},...,v_{k}\}$ where $k\geq 1$ such that $\deg(v_{i})\geq 4$ for $v_{i}\in I$ and $(v_{i-1},v_{i+1})$ is an edge,
  2. 2.
    $\deg(v_{i})=3$ for $v_{i}\in V-I$ and no other edges exist between two vertices in $V-I$ except $(v_{n},v_{2})$ when $k=1$, and
  3. 3.
    if two edges connecting $I$ to $V-I$ intersect in the interior of the polygon, then the two vertices in $I$ are connected by an edge of the outer polygon.
Theorem A.1.
Let $G$ be a braced polygon circuit. Then $G$ is minimally 3-connected if and only if $G$ satisfies the unique interval property.
For the proof, we introduce the idea of a dual polygon for minimally 3-connected braced polygons.
Definition A.2.
For a brace $e$ in a minimally 3-connected braced polygon graph $G=(V,E)$, a dual brace $e^{\prime}$ is a pair of vertices that disconnect $(V,E-\{e\})$. A dual polygon graph $G^{*}$ of $G$ is the polygon with each brace $e$ of $G$ replaced by a dual brace $e^{\prime}$.
Observe that in order for the dual brace to disconnect $(V,E-\{e\})$, each dual brace crosses exactly one brace. Therefore, no two braces can share the same dual brace. Since there are equal numbers of braces and dual braces, each brace crosses exactly one dual brace. Note also that for any given brace, there may be more than one way to choose a dual brace. An example of a convex braced polygon and its dual is given in figure 23.
Figure 23: A convex braced polygon in blue with the dual braces in red
To prove Theorem 3.7, we first need some properties of a dual polygon.
Proposition A.2.
Let $G$ be a minimally 3-connected braced polygon circuit with $n$ vertices. For every connected segment of $k<n$ ($k>1$) vertices in the outer polygon, there can be at most $k-2$ dual braces in the induced subgraph on this segment of a dual polygon $G^{*}$.
Proof. We proceed by induction. The case for $k=2$ is trivial. Suppose that the proposition is true up to $k-1$ vertices, we prove it for $k$.
Consider the segment $\{1,2,...,k\}$. Since $G$ is 3 vertex connected, there exists a brace $(\alpha,\beta)$ such that $\beta\in\{2,...,k-1\}$ and $\alpha\in\{k+1,...,n\}$. Considering the two subsegments $\{1,...,\beta\}$ and $\{\beta,...,k\}$, they can have at most $\beta-2$ and $k-\beta-1$ dual braces in the corresponding subsegments according to the induction hypothesis. At most one dual brace can cross $(\alpha,\beta)$. Therefore, we can have at most
$(\beta-2)+(k-\beta-1)+1=k-2$
dual braces on the segment $\{1,...,k\}$. ∎
Proposition A.3.
Let $G$ be a minimally 3-connected braced polygon circuit with $n$ vertices. Each dual polygon $G^{*}$ of $G$ is 3 vertex connected.
Proof. We proceed by contradiction. Suppose that the proposition is false and $\{1,k\}$ disconnects $G^{*}$. In other words, no dual brace can cross $\{1,k\}$ and all dual braces are distributed within two subsegments.
First, observe that $(1,k)$ cannot be a dual brace itself. Otherwise, the segment $\{1,2,...,k\}$ has at most $k-2$ dual braces, and the segment $\{k,k+1,...,1\}$ has at most $n-k$ dual braces by proposition A.2. Since $(1,k)$ exists in both segments, there are at most $n-3$ dual braces in $G^{*}$, contradicting $G$ being a circuit.
Next, each induced subgraph of $G^{*}$ on $\{1,2,...,k\}$ and $\{k,k+1,...,1\}$ must be 2 vertex connected. Otherwise, some vertex further splits a segment into 2 subsegments. By proposition A.2, that segment cannot have maximum number of dual braces. However, we need both segments to have the maximum number of dual braces, i.e. $k-2$ and $n-k$, so that the dual graph must have $n-2$ dual braces as proposed.
Finally, since $G$ is 3 vertex connected, there exists a brace $e$ from $\{2,...,k-1\}$ to $\{k+1,k+2,...,n\}$. However, since both $\{1,2,...,k\}$ and $\{k,k+1,...,1\}$ are 2 vertex connected, $e$ must cross two dual braces. This gives us a contradiction. ∎
Proposition A.4.
Let $G$ be a minimally 3-connected braced polygon circuit with $n$ vertices. All degree 3 vertices of $G$ are consecutive on the outer polygon.
Proof. Suppose that we have a segment of $G$, $\{k+1,...,n\}$, with all vertices having degree 3. If $k=1$, then the degree 3 vertices are trivially consecutive. We consider $k>1$. If every vertex of the segment is connected to $\{1,...,k\}$ through a brace, then there are at most $n-k$ braces with one end in the segment. Therefore, the induced subgraph of $G$ on $\{1,...,k\}$ has the maximal number: $k-2=(n-2)-(n-k)$ braces. For each brace that has both ends in $\{k+1,...,n\}$, the number of braces in $\{1,...,k\}$ increases by one. Therefore, there cannot be any internal brace in the segment $\{k+1,...,n\}$ in order for $\{1,...,k\}$ to have at most $k-2$ braces.
Suppose that there is a segment of $G$, $\{k+1,...,n\}$, of maximal length with every vertex being degree 3, we can assume vertices $1,k$ have degree at least 4 and that the segment $\{1,...,k\}$ has $k-2$ braces in $G$. We prove the following: if the induced subgraph of $G$ on a segment $\{1,...,k\}$ has $k-2$ braces, and if both $1$ and $k$ has degree at least 4, then each vertex $12,...,k-$ 1 must have degree at least 4.
We proceed by induction on $k$. The claim is trivial for $k=2$. Suppose that the claim is true for segments with length $11,...,k-$ 1, we show it for length $k$.
Since $G^{*}$ is 3-vertex connected by proposition A.3, there exists a dual brace $(\alpha,\beta)$ such that $\alpha\in\{k+1,...,n\}$ and $\beta\in\{2,...,k-1\}$. Since only one brace can cross $(\alpha,\beta)$, each subsegment $\{1,...,\beta\}$ and $\{\beta,...,k\}$ must have maximum number of braces, $\beta-2$ and $k-\beta-1$ respectively. If $\beta$ has degree at least $4$, then the proof is finished by the induction hypothesis.
If $\beta$ has degree $3$, then $(\beta-1,\beta+1)$ can be a dual brace. This is demonstrated in Figure 24. There are two cases: $\beta\in\{3,...,k-2\}$ or $\beta\in\{2,k-1\}$. In either case, we show a contradiction.
Figure 24: The braces are blue and the dual braces are red. If $\beta$ is in the middle of the segment $(1,...,k)$, the segment needs 3 more braces but it can only have 2 more crossing the two dual braces. If $\beta=k-1$, then the brace connected to $k$ must stay inside $\{1,...,k\}$. In this case, $(\alpha,k)$ will separate the graph
Case 1: If $\beta\in\{3,...,k-2\}$, then $\{1,...,\beta-1\}$, $\{\beta-1,...,\beta\}$, $\{\beta,...,\beta+1\}$, and $\{\beta+1,...,k\}$ are four subsegments separated from each other by the two dual braces $(\alpha,\beta)$ and $(\beta-1,\beta+1)$. They can have at most $\beta-3$, 00, 00, $k-\beta-2$ braces, respectively, within each subsegment. Furthermore, $2$ more braces can have two ends in distinct subsegments by crossing a dual brace. This gives only $(\beta-3)+(k-\beta-2)+2=k-3$ braces on the segment $\{1,...,k\}$, contradicting our assumption that $\{1,...,k\}$ has $k-2$ braces. This case is on the left side of Figure 24.
Case 2: If $\beta=2$ or $\beta=k-1$, without loss of generality, we can assume $\beta=k-1$. By our assumption of $k$ having degree at least 4, $\alpha\not=k+1$. We proceed to show the contradiction that $G$ cannot be $3$ vertex connected. The segment $\{1,...,k-1\}$ has exactly $k-3$ braces because only one brace can cross $(\alpha,\beta)$. In order for $\{1,...,k\}$ to have $k-2$ braces, the brace whose dual is $(\alpha,\beta)$ must be in the segment $\{1,...,k\}$. If $k\geq 4$, then the brace with one end $k-1$ has its other end in $\{1,...,k-2\}$ in order for the number of braces in $\{1,...,k-1\}$ to be $k-3$. In this case, $(\alpha,k)$ separates the graph as demonstrated by the right side of Figure 24. If $k=3$, $(1,3)$ is the only possible brace with dual $(\alpha,\beta)$. Furthermore, $\alpha\not=4,n$ when $k=3$ because vertex $1$ has degree at least 4 by our assumption. As a result, either $(\alpha,3)$ or $(\alpha,1)$ will separate the graph, depending on whether the brace with one end $2$ has the other end in $\{4,5,...,\alpha\}$ or $\{\alpha+1,\alpha+2,...,n\}$. ∎
Proposition A.5.
Let $G$ be a minimally 3 vertex connected braced polygon circuit with $n$ vertices. If the segment $\{n,1,2,...,k,k+1\}$ has $k$ braces and vertices $1,...,k$ all have degree at least 4 in $G$, then $(n,2),(1,3),...,(k-1,k+1)$ are braces of $G$.
Proof. We proceed by induction on $k$. The claim is trivial for $k=1$. Suppose that the proposition is true for $11,...,k-$ 1, we prove it for $k$.
Since the $G^{*}$ is 3-vertex connected by proposition A.3, there exists a dual brace $(\alpha,\beta)$ such that $\alpha\in\{k+2,...,n-1\}$ and $\beta\in\{1,...,k\}$. Because only one brace can cross $(\alpha,\beta)$, each smaller segment $\{n,1,...,\beta\}$ and $\{\beta,...,k,k+1\}$ must have maximum number of braces in $G$.
If $\beta\in\{2,...,k-1\}$, by induction hypothesis, $(n,2),...,(\beta-2,\beta)$ and $(\beta,\beta+2),...,(k-1,k+1)$ are braces. It remains to show $(\beta-1,\beta+1)$ is a brace. We have used $k-1$ braces, so there must be only one remaining brace on the segment $\{n,1,...,k,k+1\}$. Since both subsegments $\{n,1,...,\beta\}$ and $\{\beta,...,k,k+1\}$ have the maximum number of braces, the remaining one must cross the dual brace $(\alpha,\beta)$.
Figure 25: The dual braces cannot have their left ends in either $k+1$ or $n$. Otherwise, either $1$ or $k$ cannot have degree at least 4
Consider the dual braces of $(\beta,\beta-2)$ and $(\beta,\beta+2)$, they must have one end in $\{\beta-1,\beta+1\}$ and the other end in $\{k+1,...,n\}$. If both of them have the other end being either $k+1$ or $n$, then $(k+1,n)$ is a valid candidate for the last brace in $\{1,...,k\}$, otherwise $(\beta-1,\beta+1)$ is the only choice. The earlier case cannot occur because both $1$ and $k$ have degree at least $4$. This is shown on the left of Figure 25.
If $\beta=1$ or $\beta=k$, without loss of generality we assume $\beta=k$, by the induction hypothesis, $(n,2),...,(k-2,k)$ are braces. It remains to show $(k-1,k+1)$ is a brace. By counting we can have only one more brace and one end of it must be $k+1$. Consider the dual brace of $(k-2,k)$. One end of it must be $k-1$, and the other end can only be one of $k+2,...,n-1$ so that vertex $1$ and $k$ can have degree at least $4$. This is shown on the right of Figure 25. In this case, $(k-1,k+1)$ must be a brace. ∎
We are now ready to prove Theorem A.1.
Proof of Theorem A.1. Suppose that $G$ is minimally 3-connected. Notice that Propositions A.4 and A.5 give the first two parts of the unique interval property. To show the last property, suppose that the two lines $(v_{i},v_{k})$ and $(v_{j},v_{l})$ intersect in the interior, $v_{i},v_{j}\in I$ and $v_{k},v_{l}\in V-I$. Without loss of generality, we assume $i<j$. If $|j-i|\geq 2$, then deleting edge $(i,i+2)$ (guaranteed to exist by Propositions A.4 and A.5) will not affect 3-connectivity. To see this, deleting the edge $(i,i+2)$ only exposes $v_{i+1}$ to the interior, but no dual brace from $v_{i+1}$ can separate the graph by connecting $v_{i+1}$ to $V-I$ without crossing $(v_{i},v_{k})$ or $(v_{j},v_{l})$. Hence, $G$ satisfies the unique interval property.
Conversely, suppose that $G$ satisfies the unique interval property with respect to an interval $I$. Since each vertex in $V-I$ has degree three, $G-e$ is not 3-connected if an edge $e$ is incident to a vertex in $V-I$. Suppose that an edge $e=(i,i+2)$ is deleted with $i,i+2\in I$. Then the last condition of the universal interval property guarantees that a line through $i+1$ can separate the graph. Hence $G$ is minimally 3-connected. ∎